# Minimax and adaptive estimation of general linear functionals under sparsity

### v1

@@ 3 Adaptation to unknown sparsity
Equation 11 differs from Equation 5 (for known sparsity) because its right-hand side involves the additional term $\sqrt{\log(es)}$ in the denominator. This extra factor is crucial for achieving adaptivity. The existence and uniqueness of $\lambda_{*}(s)$ can be ensured by Lemma 11 in the appendix.
@@
We now introduce the adaptive estimator that does not rely on the knowledge about $\|\theta\|_{0}$. Motivated by [10], the adaptive estimator is selected from the collection of nonadaptive estimators $\{\hat{L}_{s}^{*}\}$ via a Lepski type scheme. Define $s_{*}=\max\left\{s\in[d]:\lambda_{*}(s)>0\right\}$ and $s_{0}=s_{*}+1$. For any $s\in[1,d]$, define
@@
The following result shows that for any $\eta$, $\Phi_{\rm adp}(s;\eta)$ is “almost increasing” in $s$ while $\Phi_{\rm adp}(s;\eta)/(s^{2}\log(es))$ is “almost decreasing” in $s$.
@@
For any $\alpha>0$ and $\gamma\in(0,2)$, there exist some constants $\tau_{\alpha},C_{0},C_{1},C_{2}>0$ such that the following holds for all $\tau\geq\tau_{\alpha}$. If an estimator $\hat{T}$ satisfies that
@@
Motivated by the illustration following Theorem 4, we consider $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the candidate adaptive rate for general loading vectors $\eta$. In the homogeneous loading case in [10], both the minimax rate and the lower bound precluding any local improvement are available in closed form, which readily identifies the adaptive rate. However, such explicit formulas are typically unavailable for general $\eta$, so it is not evident that an estimator with local improvement must incur a larger loss on $\Theta_{1}$. To establish $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the adaptive rate, we consider the following assumption on $\eta$.
Assumption 1.
There exist some positive constants $c$ and $\gamma_{0}\in(0,2)$, and some diverging $s_{\rm cut}\in[d]$ as $d\to\infty$ such that
@@
In the above assumption, condition (25) guarantees that the candidate adaptive rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ is of the same order as the minimax (nonadaptive) rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ when the sparsity level is not too large ($s\leq s_{\rm cut}$), while condition (26) essentially ensures that the ratio $\Phi_{\rm adp}(s;\eta)/\Phi_{\rm o}(1;\eta)$ grows faster than some polynomial in $s$ (with exponent $\gamma_{0}$) for large sparsity levels ($s\geq s_{\rm cut}$). In Section 5, we will show that Assumption 1 is mild and is satisfied by many loading vectors, including the homogeneous loadings.
@@
    • For $s\geq s_{\rm cut}$, the rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ remains near-optimal in the sense that, by Proposition 2, it exceeds the minimax (nonadaptive) rate by at most a logarithmic factor $\log^{2}(es)$. This means, for any function $\psi^{\prime}(s;\eta)$ being the rate of some estimator in the sense of Equation 22, it holds that $\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\gtrsim\log^{-2}(es)$. Suppose $\min_{s\geq s_{\rm cut}}\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\to 0$ as $d\to\infty$ (cf. Equation 23), then there is a diverging sequence of $\tilde{s}$ and a sequence of $\hat{T}$ such that Equation 19 in Theorem 4 holds for $s=\tilde{s}$. Theorem 4 (with $\gamma=\gamma_{0}/2$) then establishes that the maximal risk of $\hat{T}$ over the class $\Theta_{1}$ must be lower bounded by $\sigma^{2}C_{2}\Phi_{\rm adp}(\tilde{s};\eta)/{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. It then follows from Equation 26 in Assumption 1 that $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\gtrsim{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. Since $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\min_{s\geq s_{0}}[\psi^{\prime}(s)/\Phi_{\rm adp}(s;\eta)]\gtrsim(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}\log^{-2}(e(s\wedge s_{0}))\to\infty$, Equation 24 in Definition 2 is met with $\bar{s}=1$.
@@ 4 General exponentially decaying noise distributions
In this section, we study the nonadaptive estimation problem when the noise distribution has sub-Weibull tails but is not necessarily symmetric (referred to as non-symmetric).
@@
We denote by $\mathcal{H}_{\alpha,\tau}^{\otimes}$ the class of product distributions on $\mathbb{R}^{d}$ whose marginals belong to $\mathcal{H}_{\alpha,\tau}$. In other words, $\mathcal{H}_{\alpha,\tau}^{\otimes}=\{\otimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{H}_{\alpha,\tau},~~j=1,\ldots,d\}$. Our goal is to characterize the minimax optimal rate for estimating the linear functional $L(\theta)$ over $\theta\in\Theta_{s}$ and $\mathcal{P}_{\xi}\in\mathcal{H}_{\alpha,\tau}^{\otimes}$ for any $s\in[d]$. In particular, we establish nonasymptotic upper and lower bounds on the minimax risk that match up to at most logarithmic factors.
For any integer $s\in[d]$, define
and consider the estimator
where $c_{\mathcal{H}}>0$ is a constant depending only on $\alpha$ and $\tau$.
The following theorem gives an upper bound on the risk of the estimator $\hat{L}_{\cal H}$.
Theorem 5.
We now turn to establishing a lower bound on the minimax risk. To this end, for any $s\in[d]$, let $\lambda_{\cal H}\geq 0$ be the unique solution to
@@
The next theorem provides a lower bound on the maximal risk of any estimator of $L(\theta)$ when the noise distribution belongs to $\mathcal{H}_{\alpha,\tau}^{\otimes}$.
@@
Theorem 6 establishes two complementary lower bounds. The bound in (31) follows directly from Theorem 1, since ${\cal G}_{\alpha,\tau}^{\otimes}\subseteq{\cal H}_{\alpha,\tau}^{\otimes}$. We refer to this as the symmetric lower bound, as it is the minimax rate under symmetric noise distributions.
@@
This proposition shows that the symmetric lower bound in (31) is optimal up to a logarithmic factor, matching the upper bound in Theorem 5. Moreover, the risk of the pure plug-in estimator $\sum_{j=1}^{d}\eta_{j}y_{j}$ is bounded by $\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}$. Combining this with the symmetric lower bound, Proposition 4 implies that once $s\gtrsim\sqrt{d}$, the minimax rate is given by $\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}$ and becomes independent of the sparsity level $s$.
In contrast, the bound in (33) arises from a construction that exploits the asymmetry of the noise distribution; we therefore call it the asymmetric lower bound. In Section 5, we show that under the homogeneous loading vector, the asymmetric lower bound can be strictly larger than the symmetric one for $s=o(\sqrt{d})$. The key idea is to use the Hellinger distance, rather than the $\chi^{2}$-divergence, to control the total variation distance. This approach has appeared in [10, 24], though primarily in the context of estimating the quadratic functional $\|\theta\|_{2}^{2}$, and thus cannot be directly applied here. While the construction technique is of independent interest, we omit further details due to space limitations.
@@ 5.2 Two-phase loading vector
For convenience, we parameterize $d_{0}$ and $\lambda$ as $d_{0}=d^{\gamma_{d}}$ and $\lambda=d^{\gamma_{\lambda}}$ for some constants $\gamma_{d}>0$ and $\gamma_{\lambda}>0$. To uncover an interesting phase transition phenomenon as discussed below, we assume $2\gamma_{\lambda}+\gamma_{d}<1$. We summarize the rates in Table 1.
$\Phi_{\rm o}(s;\eta)$ $\Phi_{\rm adp}(s;\eta)$
$s\lesssim\frac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d}$ $d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\gamma_{d}\alpha/2}}{s^{\alpha}}\right)$ $\begin{array}[]{l}\displaystyle s^{2}\!\left[d^{2\gamma_{\lambda}}\log\!\left(1+\frac{d^{\gamma_{d}\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)+\right.\\ \left.\qquad\log\!\left(1+\frac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)\right]\end{array}$
$s\gtrsim\frac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d}$ $s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right)$
Table 1: Optimal rates in the two-phase case: minimax rates $\Phi_{\rm o}$ and adaptive rates $\Phi_{\rm adp}$.
@@
It is worth mentioning that the transition in the minimax rate from $\Phi_{\rm o}(s;\lambda\mathbf{1}_{d_{0}})$ to $\Phi_{\rm o}(s;\mathbf{1}_{d})$ is not observed in previous studies on homogeneous loadings.
@@ 5.3 Exponentially Decaying Loading Vector
We allow the function $\phi$ to depend on $d$; for example, $\phi(x)=2(x/d)^{\gamma}$ for $\gamma\geq 1$. Given $\phi$, define $j_{0}=\min\{\,j\mid\eta_{j}<1/2\,\}$ if this minimum is taken over a non-empty set, and $j_{0}=d+1$ otherwise. Then the optimal rates are given by
@@ 6 Extensions
6.1 Estimation with unknown noise level
The estimation problem with an unknown noise level has been widely studied in the literature; see, for example, [9, 10, 11, 7]. A common approach is to first construct an estimator of $\sigma$ and then substitute it into the original procedure in place of the true noise level. For example, [11] proposed a median-of-means estimator:
Let $\gamma_{\alpha,\tau}\in(0,1/2]$ be a constant chosen sufficiently small and only related to $\alpha$ and $\tau$. Divide $\{1,\ldots,d\}$ into $m=\lfloor\gamma_{\alpha,\tau}d\rfloor$ disjoint subsets $B_{1},\ldots,B_{m}$, each of cardinality
@@ 6.1 Estimation with unknown noise level
$\hat{\sigma}^{2}={\rm median}(\bar{\sigma}_{1}^{2},\ldots,\bar{\sigma}_{m}^{2}),\qquad\text{ where }\bar{\sigma}_{i}^{2}=\frac{1}{|B_{i}|}\sum_{j\in B_{i}}y_{j}^{2},\quad i=1,\ldots,m.$ (36)
@@
Based on the estimator $\hat{\sigma}$ in (36), most of our results and phenomena for estimating the linear functional $L(\theta)$ extend naturally to the setting with an unknown noise level. Due to space limitations, we restrict attention to the estimation problem with known sparsity. Replacing $\sigma$ in the estimator $\hat{L}_{s}$ defined in (7) with $\hat{\sigma}$ defined in (36) and adjusting the constant in front of the threshold, we obtain the estimator
@@
For any $\alpha,\tau>0$, there are some constants $\gamma_{\alpha,\tau}$ and $C$ depending only on $\alpha$ and $\tau$ such that if $1\leq s<\lfloor\gamma_{\alpha,\tau}d\rfloor/4$, then

### v2

@@ 3 Adaptation to unknown sparsity
Equation 11 differs from Equation 5 (for known sparsity) because its right-hand side involves the additional term $\sqrt{\log(es)}$ in the denominator. This extra factor is crucial for achieving adaptivity. The existence and uniqueness of $\lambda_{*}(s)$ can be ensured by Lemma 11 in the supplementary material.
@@
We now introduce the adaptive estimator that does not rely on knowledge about $\|\theta\|_{0}$. Motivated by [10], the adaptive estimator is selected from the collection of nonadaptive estimators $\{\hat{L}_{s}^{*}\}$ via a Lepski type scheme. Define $s_{*}=\max\left\{s\in[d]:\lambda_{*}(s)>0\right\}$ and $s_{0}=s_{*}+1$. For any $s\in[1,d]$, define
@@
The following result shows that for any $\eta$, $\Phi_{\rm adp}(s;\eta)$ is “almost increasing” in $s$ while the ratio $\Phi_{\rm adp}(s;\eta)/(s^{2}\log(es))$ is “almost decreasing” in $s$.
@@
For any $\alpha>0$ and $\gamma\in(0,2)$, there exist constants $\tau_{\alpha},C_{0},C_{1},C_{2}>0$ such that the following holds for all $\tau\geq\tau_{\alpha}$. If an estimator $\hat{T}$ satisfies that
@@
Motivated by the illustration following Theorem 4, we consider $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the candidate adaptive rate for general loading vectors $\eta$. In the homogeneous loading case in [10], both the minimax rate and the lower bound precluding any local improvement are available in closed form, which readily identifies the adaptive rate. However, such explicit formulas are typically unavailable for general $\eta$, so it is not evident that an estimator with local improvement must incur a larger loss on $\Theta_{1}$. To establish $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the adaptive rate, we consider the following condition on $\eta$.
Condition 1.
There exist constants $c>0$ and $\gamma_{0}\in(0,2)$, and a sequence $s_{\rm cut}=s_{\rm cut}(d)\in[d]$ that diverges as $d\to\infty$ such that
@@
In the above condition, Equation 25 guarantees that the candidate adaptive rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ is of the same order as the minimax (nonadaptive) rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ when the sparsity level is not too large ($s\leq s_{\rm cut}$), while condition (26) essentially ensures that the ratio $\Phi_{\rm adp}(s;\eta)/\Phi_{\rm o}(1;\eta)$ grows faster than some polynomial in $s$ (with exponent $\gamma_{0}$) for large sparsity levels ($s\geq s_{\rm cut}$).
1 only depends on the loading vector $\eta$ and is thus verifiable. In Section 5, we will show that 1 is mild and is satisfied by many loading vectors, including the homogeneous loadings.
@@
    • For $s\geq s_{\rm cut}$, the rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ remains near-optimal in the sense that, by Proposition 2, it exceeds the minimax (nonadaptive) rate by at most a logarithmic factor $\log^{2}(es)$. This means that for any function $\psi^{\prime}(s;\eta)$ being the risk rate of some estimator (in the sense of Equation 22), we have $\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\gtrsim\log^{-2}(es)$. Suppose $\min_{s\geq s_{\rm cut}}\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\to 0$ as $d\to\infty$ (cf. Equation 23), then there is a diverging sequence of $\tilde{s}$ and a sequence of $\hat{T}$ such that Equation 19 in Theorem 4 holds for $s=\tilde{s}$. Theorem 4 (with $\gamma=\gamma_{0}/2$) then establishes that the maximal risk of $\hat{T}$ over the class $\Theta_{1}$ must be lower bounded by $\sigma^{2}C_{2}\Phi_{\rm adp}(\tilde{s};\eta)/{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. It then follows from Equation 26 in Assumption 1 that $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\gtrsim{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. Since $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\min_{s\geq s_{0}}[\psi^{\prime}(s)/\Phi_{\rm adp}(s;\eta)]\gtrsim(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}\log^{-2}(e(s\wedge s_{0}))\to\infty$, Equation 24 in Definition 2 is met with $\bar{s}=1$.
@@ 4 General exponentially decaying noise distributions
In the previous sections, we have focused on symmetric noise distributions. In this section, we show that the symmetry assumption plays a crucial role in establishing the preceding results; once this assumption is relaxed, the minimax rate may increase.
To formalize the problem, we consider the following larger class of noise distributions that are not necessarily symmetric (referred to as non-symmetric).
@@
We denote by $\mathcal{H}_{\alpha,\tau}^{\otimes}$ the class of product distributions on $\mathbb{R}^{d}$ whose marginal distributions belong to $\mathcal{H}_{\alpha,\tau}$. In other words, $\mathcal{H}_{\alpha,\tau}^{\otimes}=\left\{\bigotimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{H}_{\alpha,\tau},\;j=1,\ldots,d\right\}$.
For any $s\in[d]$, let $\lambda_{\cal H}\geq 0$ be the unique solution to
@@
The following theorem provides a lower bound on the maximal risk of any estimator of $L(\theta)$ when the noise distribution belongs to $\mathcal{H}_{\alpha,\tau}^{\otimes}$.
@@
Theorem 5 establishes two complementary lower bounds. The bound in (28) follows directly from Theorem 1, since ${\cal G}_{\alpha,\tau}^{\otimes}\subseteq{\cal H}_{\alpha,\tau}^{\otimes}$ according to Definitions 1 and 3. We refer to this as the symmetric lower bound, as it is the minimax rate under symmetric noise distributions.
In contrast, the bound in (30) is obtained from a construction that explicitly exploits the asymmetry of the noise distribution; we therefore refer to it as the asymmetric lower bound. As demonstrated in Section 5, when the loading vector is homogeneous, the asymmetric lower bound can be strictly larger than its symmetric counterpart for $s=o(\sqrt{d})$, indicating that the results established in the symmetric setting no longer hold once the noise distribution loses symmetry.
The key technical ingredient of the proof for (30) is the use of the Hellinger distance, rather than the $\chi^{2}$-divergence, to control the total variation distance. This idea has appeared in earlier works such as [10] and [23], though primarily in the context of estimating the quadratic functional $\|\theta\|_{2}^{2}=\sum_{j=1}^{d}\theta_{j}^{2}$, and hence cannot be directly applied to our setting. To overcome this challenge, we again employ the random sparsity strategy and construct a novel prior. The proof may be of independent theoretical interest; detailed derivations are omitted here for brevity, and the full arguments are provided in the supplementary material.
The lower bound highlights the challenge introduced by asymmetry in the noise distribution. To complement this result, we establish an upper bound that generally matches the lower bound up to a logarithmic factor and can coincide with it in certain regimes. For any integer $s\in[d]$, define
and consider the estimator
where $c_{\mathcal{H}}>0$ is a constant depending only on $\alpha$ and $\tau$.
The next theorem gives an upper bound on the risk of the estimator $\hat{L}_{\cal H}$.
Theorem 6.
The upper bound established in Theorem 6 is nearly sharp. Section 5.1 shows an example where this upper bound can match the lower bounds in Theorem 5 in some regimes. In general, we have the following proposition for comparing this upper bound with the minimax rate characterized in Theorem 1.
@@
This proposition indicates that the cost of non-symmetry is at most a logarithmic factor. Furthermore, the symmetric lower bound $\sigma^{2}\Phi_{\mathrm{o}}(s;\eta)$ matches the upper bound in Theorem 3 up to logarithmic terms. In particular, in the dense regime where $s\gtrsim\sqrt{d}$, both the symmetric lower bound and the upper bound are of order $\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}$, yielding the exact minimax rate. In this regime, the rate is independent of both the sparsity level $s$ and the symmetry of the noise distribution.
@@ 5.2 Two-phase loading vector
For convenience, we parameterize $d_{0}$ and $\lambda$ as $d_{0}=d^{\gamma_{d}}$ and $\lambda=d^{\gamma_{\lambda}}$ for some constants $\gamma_{d}>0$ and $\gamma_{\lambda}>0$. To uncover an interesting phase transition phenomenon as discussed below, we assume $2\gamma_{\lambda}+\gamma_{d}<1$.
The minimax rate is given by
$\Phi_{\mathrm{o}}(s;\eta)$ $\asymp\begin{cases}d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\dfrac{d^{\gamma_{d}\alpha/2}}{s^{\alpha}}\right),&s\lesssim\dfrac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d},\\[6.0pt] s^{2}\log^{2/\alpha}\!\left(1+\dfrac{d^{\alpha/2}}{s^{\alpha}}\right),&s\gtrsim\dfrac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d},\end{cases}$
while the adaptive rate is given by
$\Phi_{\mathrm{adp}}(s;\eta)$ $\asymp s^{2}\!\left[d^{2\gamma_{\lambda}}\log\!\left(1+\dfrac{d^{\gamma_{d}\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)+\log\!\left(1+\dfrac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)\right].$
@@ 5.3 Exponentially Decaying Loading Vector
We allow the function $\phi$ to depend on $d$; for example, $\phi(x)=2(x/d)^{\gamma}$ for $\gamma\geq 1$. Given $\phi$, define $j_{0}=\min\{\,j\mid\eta_{j}<1/2\,\}$ if the set is nonempty; otherwise set $j_{0}=d+1$. The optimal rates are given by
@@ 6 Extensions
6.1 Estimation with unknown noise variance
The estimation problem with an unknown noise variance has been widely studied in the literature; see, for example, [9, 10, 11, 7]. A common approach is to first construct an estimator of $\sigma^{2}$ and then substitute it into the original procedure in place of the true noise variance. For example, [11] proposed a median-of-means estimator:
Let $\gamma_{\alpha,\tau}\in(0,1/2]$ be a constant chosen sufficiently small, depending only on $\alpha$ and $\tau$. Divide $\{1,\ldots,d\}$ into $m=\lfloor\gamma_{\alpha,\tau}d\rfloor$ disjoint subsets $B_{1},\ldots,B_{m}$, each of cardinality
@@ 6.1 Estimation with unknown noise variance
$\hat{\sigma}^{2}=\operatorname{median}(\bar{\sigma}_{1}^{2},\ldots,\bar{\sigma}_{m}^{2}),\qquad\text{ where }\bar{\sigma}_{i}^{2}=\frac{1}{|B_{i}|}\sum_{j\in B_{i}}y_{j}^{2},\quad i=1,\ldots,m.$ (36)
@@
Based on the estimator $\hat{\sigma}^{2}$ in (36), most of our results for estimating the linear functional $L(\theta)$ extend naturally to the setting with an unknown noise variance. Due to space limitations, we restrict attention to the estimation problem with known sparsity. Replacing $\sigma$ in the estimator $\hat{L}_{s}$ defined in (7) with $\hat{\sigma}$ defined in (36) and adjusting the constant in front of the threshold, we obtain the estimator
@@
For any $\alpha,\tau>0$, there exist constants $\gamma_{\alpha,\tau}$ and $C$ depending only on $\alpha$ and $\tau$, such that if $1\leq s<\lfloor\gamma_{\alpha,\tau}d\rfloor/4$, then